Circular No. 8566

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/1981 W1 (SOLWIND) R. Kracht, Elmshorn, Germany, reports that he has found a previously

unrecognized Kreutz sungrazing comet on Solwind spacecraft images taken on 1981 Nov. 20. C/1981 W1, which showed a small tail, was fainter than the previous Solwind comets (cf. IAUC 4129, 4229, 4230) — perhaps of mag 2–3. The observations below were provided by Kracht.

1981 UT  α_{2000}  δ_{2000}
Nov. 20.059  15^{h}37.^{m}9  −21^{o}52′
20.062  15 37.9  −21 52
20.066  15 38.0  −21 51
20.069  15 38.0  −21 50
20.073  15 38.1  −21 49
20.076  15 38.2  −21 48
20.079  15 38.2  −21 49
20.083  15 38.3  −21 47
20.086  15 38.3  −21 47
20.090  15 38.4  −21 46
20.265  15 41.3  −21 12
20.272  15 41.4  −21 11
20.275  15 41.5  −21 09
20.279  15 41.6  −21 09

The following parabolic orbital elements are by B. G. Marsden:

T = 1981 Nov. 20.61 TT  ω = 97.^{o}36
Ω = 24.63 }2000.0
q = 0.0048 AU  i = 135.48

COMET C/2005 K2 (LINEAR) Additional offsets of the secondary component from the primary (cf.

IAUC 8548) from CCD images, reported via E. Jehin, European Southern Observatory: June 14.932 UT, 40′′.0 in p.a. 47^{o}.1 (O. Nielsen, Rijswijk, The Netherlands, 0.20-m reflector); 14.940, 40′′.2 in p.a. 47^{o}.7 (E. Bryssinck, Antwerp, Belgium, 0.28-m reflector).

Visual total-magnitude estimates: June 11.87 UT, 9.0 (G. Santa, Szeged, Hungary, 0.11-m reflector); 13.93, 9.6 (A. Baransky, Bucha, Ukraine, 0.20-m reflector); 17.91, 9.2 (J. J. Gonzalez, Leon, Spain, 0.20-m reflector); 24.35, 9.9 (D. A. J. Seargent, Cowra, N.S.W., 0.25-m reflector).

2005 July 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green